# X-Vine Models for Multivariate Extremes

### v1

@@ 9 Application: US flight delay data
We apply X-vine models to investigate extremal dependence among large flight delays in the US flight network. The raw data-set is accessible through the US Bureau of Transportation Statistics [^1] https://www.bts.dot.gov. These flight delay data were analysed by Hentschel et al. (2022) who first selected airports with a minimum of 1 000 flights per year and applied a $k$-medoids clustering algorithm to identify homogeneous clusters in terms of extremal dependence. This clustering approach not only makes the analysis suitable for modelling extremal dependence but also reveals shared frequent flight connections between airports and similar geographical characteristics in each cluster. Focusing on the Hüsler–Reiss family, Hentschel et al. (2022) fitted an extremal graphical model to large flight delays of airports in the Texas cluster to investigate conditional independence.
For the purpose of model comparison, we also analyse daily total delays (in minutes) in the Texas cluster. The cluster comprises $d=29$ airports and counts $n=3603$ days from 2005 to 2020, during which all airports have recorded measurements. This pre-processed data is available through the R-package graphicalExtremes (Engelke et al., 2022). A graphical representation of the actual flight connections between airports is presented in Fig. LABEL:fig:FlightGraphDomain. While the graph derived from flight connections does not guarantee the optimal graph in terms of extremal behavior, it offers valuable insights into the flight network.
@@
Let $\widehat{\bm{x}}_{i},$ $i=1,\ldots,n,$ denote the observed measurements. Following the standardisation process with the rank transformation as in Section 7.2, we take sub-samples $\widehat{\bm{z}}_{i}=(n/k)\widehat{\bm{u}}_{i}$ for $i\in K$ where $k/n$ is the threshold of the uniform margin $\widehat{u}_{i,j}$. We choose $k$ such that $k/n=0.13$ in order to have a large enough effective sample size in relation to the number of variables.
@@
Assigning $\widehat{\chi}_{e}$ for $e\in E_{1}$ and $\widehat{\tau}_{e}$ for $e\in E_{j}$, $j\geqslant 2$, as edge weights, we select the vine tree structure sequentially as in Section 7.3. The first maximum spanning tree is obtained with a range of estimated tail dependence coefficients $\widehat{\chi}_{e}$ between $0.42$ and $0.65$ and a mean of $0.55$ in Fig. 9 in Appendix F.3. Out of the $d-1=28$ edges in $T_{1}$, selected bivariate tail copula families include the Hüsler–Reiss ($8$ edges), negative logistic ($3$ edges) and Dirichlet models ($17$ edges). Subsequent trees with a total of $378$ edges are then chosen with the following pair-copula families (the number of edges inside parentheses): Independence (201), Gaussian (31), Clayton (8), Gumbel (26), Frank (34), Joe (36), Survival Clayton (31), Survival Gumbel (9), and Survival Joe (2). As the tree level rises, we observe a decline in residual dependence and an increase in the number of independence copulas. The X-vine algorithm sets all pair copulas to the independence copula from tree level $21$, that is, the vine is truncated at level $q=20$.
@@
To evaluate the X-vine model’s efficacy in capturing extremal dependence for large flight delays between airports in the Texas cluster, we estimate bivariate and trivariate tail dependence coefficients. We compare empirical pairwise tail dependence coefficients $\widehat{\chi}_{a,b}$ for $a,b\in[d]$ with those from the fitted X-vine model in Fig. LABEL:fig:ChiflightFullXvine. Additionally, Fig. LABEL:fig:Chi3FlightFull in Appendix F.3 compares trivariate empirical tail dependence coefficients $\widehat{\chi}_{abc}$ for $a,b,c\in[d]$ with those obtained from the fitted X-vine model. Both assessments indicate a satisfactory fit for the large flight delays data.
@@
To further explore truncated X-vine models with a lower truncation level, we use the mBIC in Eq. (7.7). Fig. LABEL:fig:mBICFlight illustrates the mBIC-optimal truncation level of $q^{*}=7$, which corresponds to 113 dependence copulas [Gaussian (17), Clayton (3), Gumbel (23), Frank (16), Joe (25), Survival Clayton (22), Survival Gumbel (6), and Survival Joe (1)] out of the 147 total edges from $T_{2}$ to $T_{7}$. The total number of bivariate model components is thus $28+113=141$. Additionally, we investigate how the truncation level changes over the threshold range of $k/n\in(0.05,0.15)$ in Fig. LABEL:fig:mBICvsQuan in Appendix F.3. It appears that the mBIC-optimal truncation levels are not overly sensitive to the choice of the threshold. The superimposed graph of the first seven vine trees is shown in Fig. LABEL:fig:FlightGraphXvineTrunc, showing for $T_{2}$ to $T_{7}$ only the 113 out of 147 edges for which a copula other than the independence one is selected. The level by level vine tree sequence from $T_{1}$ to $T_{7}$ is presented in Fig. 9 in Appendix F.3.
@@
Returning attention to model comparison, we consider extremal graphical models with the Hüsler–Reiss distribution, another sparse model for extremes, detailed in Engelke et al. (2021) and Hentschel et al. (2022). Hentschel et al. (2022) obtained the sparse Hüsler–Reiss graphical model through a structure learning method, using the EGlearn algorithm of Engelke et al. (2021) for the sequential estimation of extremal graphical structure from data. The tuning parameter $\rho\geqslant 0$ controls sparsity, and the empirical extremal variogram matrix $\widehat{\Gamma}$ is used as edge weight for the minimum spanning tree with smaller elements in $\widehat{\Gamma}$ indicating stronger extremal dependence.
Following the vignette in Engelke et al. (2022), the data set is split in half: a training set for model fitting and a test set for tuning parameter selection. We determine the optimal tuning parameter, $\rho^{*}=0.1$, by evaluating the Hüsler–Reiss log-likelihood values across tuning parameter values $\rho$ from the test set. The resulting sparse extremal graph with 148 edges at $\rho^{*}=0.1$ is shown in Fig. LABEL:fig:FlightGraphEGlearn.
@@
We assess the goodness of fit using the entire sub-samples and compare empirical tail dependence coefficients to those obtained from the fitted graphical models: the Hüsler–Reiss extremal graphical model for the flight graph in Fig. LABEL:fig:ChiFlightDomain, the truncated X-vine model with $q^{*}=7$ in Fig. LABEL:fig:ChiFlightTruncXvine, and the extremal Hüsler–Reiss graphical model with $\rho^{*}=0.1$ in Fig. LABEL:fig:ChiFlightEGlearn, respectively.
@@
The flexibility of a regular vine structure allows the selection of various (tail) copula families with the lowest AIC. Recall from Section 4.2 that the Hüsler–Reiss model arising from the X-vine specification is constructed using bivariate Hüsler–Reiss tail copula densities in the first tree and bivariate Gaussian copulas in subsequent trees. In the truncated X-vine model, 8 out of 28 edges prefer Hüsler–Reiss models in $T_{1}$, while in trees $T_{j}$ for $j\geqslant 2$, 96 out of 113 edges favour copula families other than the Gaussian one. However, uncertainty in sequential parameter estimation tends to accumulate across tree levels. Consequently, the $\chi$-plot of the truncated X-vine model shows a better fit but more variability. In contrast, the $\chi$-plot of the extremal Hüsler–Reiss graphical model has less variability but seems to biased towards higher tail dependence.
Similarly, focusing exclusively on the Hüsler–Reiss model, Fig. 12 presents a model comparison between the extremal Hüsler–Reiss graphical model based on the flight graph, the Hüsler–Reiss model derived from the X-vine structure, and the extremal Hüsler–Reiss graphical model from the EGlearn algorithm. Additional details about the fitted models and model comparison are provided in Appendix F.3.
@@ 10 Discussion
We have opened the door to the construction and the use of extremal dependence models based on regular vine tree sequences. For (ordinary) copulas, the methodology has been extensively developed in the literature since its inception more than two decades ago. Our contribution is to deliver the theoretical and methodological advances required to apply vine machinery in the extreme value context too. The key consists of a version of Sklar’s theorem applied to tail copula densities, which stand in one-to-one relation to exponent measure densities, together with a telescoping product formula for regular vines. For many popular parametric families of tail dependence models, the resulting conditional copula densities satisfy a simplifying assumption. Tail copula densities with this property, called X-vines, are specified by a triple consisting of a regular vine sequence, a vector of bivariate tail copula densities and a vector of bivariate copula densities. These bivariate building blocks can be chosen independently from another, yielding great flexibility. At the same time, the recursive properties of regular vines ensure that all computations can be reduced to operations involving the bivariate building blocks only. Further, we have studied the inverted multivariate Pareto distributions associated to these X-vines, as dependence models for excesses over high multivariate thresholds. A recursive algorithm serves to sample from such distributions, and we have developed a statistical methodology on them, in order to learn an X-vine from data. The method comprises parameter estimation, model selection, and vine structure learning, all proceeding sequentially, tree by tree. In large dimensions, truncating the vine sequence produces models with fewer parameters. We have illustrated the performance of the various methods on simulated and real data.
@@ Acknowledgments
This work is supported in part by funds from the Fonds de la Recherche Scientifique – FNRS, Belgium (grant number T.0203.21).
SUPPLEMENTARY MATERIAL
In Appendix A, we provide detailed illustrations of a number of concepts, results and methods for a particular five-dimensional regular vine sequence. Appendix B contains the proofs of all results in the paper. A converse to Sklar’s theorem in Proposition 3.1 is stated and proved in Appendix C. For readers more familiar with exponent measure densities than with tail copula densities, a number of theoretical results are reformulated in terms of exponent measure densities in Appendix D. In Appendix E, it is shown that X-vine models are the tail limits of regular vine copula models in the sense of Definition 2.1. Finally, Appendix F.2 contains additional numerical results, both for the Monte Carlo simulations in Section 8 and the US flight data case study in Section 9.
@@ Appendix A Example in dimension five
In tree $T_{3}$, consider the edge $e=\{a,b\}=\{\{\{1,2\},\{2,3\}\},\{\{2,3\},\{2,4\}\}\}\in E_{3}$, which is also a node in tree $T_{4}$. The proximity condition is satisfied since $|a\cap b|=|\{\{2,3\}\}|=1$: note that $a\cap b$ is indeed a singleton, the unique element of which is the set $\{2,3\}$. The edge $e$ has complete union $A_{e}=\{1,2,3,4\}$, which is the union of the conditioning sets $A_{a}=\left\{{1,2,3}\right\}$ and $A_{b}=\left\{{2,3,4}\right\}$. Further, $e$ has conditioning set $D_{e}=\{2,3\}=A_{a}\cap A_{b}$, and conditioned set $\mathcal{C}_{e}=A_{a}\operatorname{\bigtriangleup}A_{b}=\{1,4\}$, which is itself the union $\mathcal{C}_{e}=\mathcal{C}_{e,a}\cup\mathcal{C}_{e,b}$ of $\mathcal{C}_{e,a}=A_{a}\setminus A_{b}=\left\{{1}\right\}$ and $\mathcal{C}_{e,b}=A_{b}\setminus A_{a}=\left\{{4}\right\}$. Writing this edge in its conditioned form gives better readability: $e=(\mathcal{C}_{e,a},\mathcal{C}_{e,a};D_{e})=(14;23)$. Fig. 6 shows the regular vine sequence, where the nodes and edges have been labeled using the conditioned forms.
@@
Figure 6: A graphical representation of the five-dimensional regular vine sequence $\mathcal{V}$. The sub-vine $\mathcal{V}_{f}$ induced by $\mathcal{V}$ on the node set $A_{f}=\{2,4,5\}$ for edge $f=(25;4)$ in $T_{2}$ is shown in blue.
@@
As an illustration of Remark 2.5, consider the edge $f=\{\{2,4\},\{4,5\}\}\in E_{2}$ (or $f=(25;4)$ in its conditioned form). The vine $\mathcal{V}_{f}$ induced by $\mathcal{V}$ on the node set $A_{f}=\{2,4,5\}$ consists of two trees, $\mathcal{V}_{f}=(T_{f,1},T_{f,2})$, with edge sets $E_{f,1}=\{\{2,4\},\{4,5\}\}$ and $E_{f,2}=\{\{\{2,4\},\{4,5\}\}\}$; see Fig. 6. The product formula (2.15) yields the identity
@@
where $C_{1|3;2}(u_{1}|u_{3})=\int_{0}^{u_{1}}c_{13;2}(v,u_{3})\,\mathrm{d}v$, $R_{1|2}(x_{1}|x_{2})=\int_{0}^{x_{1}}r_{12}(t,x_{2})\,\mathrm{d}t$ and $R_{3|2}(x_{3}|x_{2})=\int_{0}^{x_{3}}r_{23}(x_{2},t)\,\mathrm{d}t$, i.e., in terms of bivariate (tail) copula densities $r_{12}$, $r_{23}$ and $c_{13;2}$ only.
@@
Proofs involving regular vine sequences often operate via a recursive argument in which a node $a\in[d]$ and the edges that contain it are peeled off, leaving a regular vine sequence on $d-1$ elements; see for instance the proof of Theorem 5.3, which relies on Lemma B.2 below. Let us illustrate the latter for the regular vine sequence $\mathcal{V}$ in Figure 6. The conditioned set of the deepest edge $e=(15;234)$ is $\mathcal{C}_{e}=\left\{{1,5}\right\}$. For each of the latter two nodes and for each level $j\in\left\{{1,\ldots,d-1}\right\}$, there is a unique edge $e_{j}\in E_{j}$ that contains this node in its complete union. For instance, for node $1$ these are the edges $(12)$, $(13;2)$, $(14;23)$ and $(15;234)$; moreover, node $1$ belongs to the conditioned set of each of these edges. Removing node $1$ and these four edges leaves a regular vine sequence on the four remaining elements $\left\{{2,\ldots,5}\right\}$, which is the sub-vine induced by $\mathcal{V}$ on the node set $\left\{{2,\ldots,5}\right\}$.
@@
Steps (2) and (4.a) leave freedom for choice, and this is why there may be multiple permutations $\sigma_{j}$ that satisfy the requirements of Lemma 6.2. For the regular vine tree $\mathcal{V}$ in Fig. 6, five possible permutations $\sigma_{j}$ for $j\in\left\{{1,\ldots,5}\right\}$ and the corresponding edge sequences $e_{j,1},\ldots,e_{j,4}$ are listed in Table 2. Another choice for $\sigma_{4}$ would have been $(4,2,5,3,1)$, for instance.
@@
Table 2: Five possible permutations $\sigma_{j}$ and the edge sequences $e_{j,1},\ldots,e_{j,4}$ for $j\in\left\{{1,\ldots,5}\right\}$ of Lemma 6.2 for the regular vine tree $\mathcal{V}$ in Figure 6.
@@
For the regular vine sequence in Fig. 6, some possible structure matrices $M$ are
@@
We illustrate the recursive nature of the sequential estimation procedure in Section 7.2 for the regular vine sequence in Fig. 6. For the edge $e=(25;4)$ in $E_{2}$, we consider pseudo-observations

### v2

@@ 9 Application: US flight delay data
We apply X-vine models to investigate extremal dependence among large flight delays in the US flight network. The raw data set is accessible through the US Bureau of Transportation Statistics [^2] https://www.bts.dot.gov. These flight delay data were analysed by Hentschel et al. (2022) who first selected airports with a minimum of 1 000 flights per year and applied a $k$-medoids clustering algorithm to identify homogeneous clusters in terms of extremal dependence. This clustering approach not only makes the analysis suitable for modelling extremal dependence but also reveals shared frequent flight connections between airports and similar geographical characteristics in each cluster. Focusing on the Hüsler–Reiss family, Hentschel et al. (2022) fitted an extremal graphical model to large flight delays of airports in the Texas cluster to investigate conditional independence.
For the purpose of model comparison, we also analyse daily total delays (in minutes) in the Texas cluster. The cluster comprises $d=29$ airports and counts $n=3603$ days from 2005 to 2020, during which all airports have recorded measurements. This pre-processed data is available through the R-package graphicalExtremes (Engelke et al., 2022). A graphical representation of the actual flight connections between airports is presented in Fig. LABEL:fig:FlightGraphDomain.
@@
Let $\widehat{\bm{x}}_{i}$, for $i=1,\ldots,n$, denote the measurements. Following the standardisation process with the rank transformation (7.1) as in Section 7.2, we take the sub-sample $\widehat{\bm{z}}_{i}=(n/k)\widehat{\bm{u}}_{i}$ for $i\in K=\bigcup_{j\in[d]}K_{j}$ as in (7.2). We choose $k=0.13n$ in order to have a large enough effective sample size with respect to the number of variables.
@@
Assigning $\widehat{\chi}_{e}$ for $e\in E_{1}$ and $\widehat{\tau}_{e}$ for $e\in E_{j}$, $j\geqslant 2$, as edge weights, we select the vine tree structure sequentially as in Section 7.3. Fig. 5(a) in the supplement shows the first maximum spanning tree. The corresponding estimated tail dependence coefficients vary between $0.42$ and $0.65$, with a mean of $0.55$. Out of the $d-1=28$ edges in $T_{1}$, selected bivariate tail copula families include the Hüsler–Reiss ($8$ edges), negative logistic ($3$ edges) and Dirichlet models ($17$ edges). Subsequent trees with a total of $378$ edges are then chosen with the following pair-copula families (number of edges in parentheses): Independence (201), Gaussian (31), Clayton (8), Gumbel (26), Frank (34), Joe (36), Survival Clayton (31), Survival Gumbel (9), and Survival Joe (2). As the tree level rises, we observe a decline in residual dependence and an increase in the number of independence copulas. The X-vine algorithm sets all pair copulas to the independence copula from $T_{21}$, that is, the vine is truncated at level $q=20$.
@@
To evaluate the X-vine model’s efficacy in capturing extremal dependence for large flight delays between airports, we compare bivariate and trivariate empirical tail dependence coefficients with those from the fitted X-vine models in Section F.2.2 in the supplement. Fig. F.6 indicates a satisfactory fit, both for the full $29$-dimensional model and for the truncated model.
@@
To further explore truncated X-vine models with a lower truncation level, we use the mBIC in Eq. (7.5). Fig. LABEL:fig:mBICFlight in the supplement illustrates the mBIC-optimal truncation level of $q^{*}=7$, which corresponds to 113 dependence copulas [Gaussian (17), Clayton (3), Gumbel (23), Frank (16), Joe (25), Survival Clayton (22), Survival Gumbel (6), and Survival Joe (1)] out of the 147 total edges from $T_{2}$ to $T_{7}$. The total number of bivariate model components is thus $28+113=141$. Additionally, we investigate how the truncation level changes over the threshold range in Fig. LABEL:fig:mBICvsQuan. It appears that the mBIC-optimal truncation levels are not overly sensitive to the choice of the threshold. The superimposed graph of the first seven vine trees is shown in Fig. LABEL:fig:FlightGraphXvineTrunc, showing for $T_{2}$ to $T_{7}$ only the 113 out of 147 edges for which a copula other than the independence one is selected. The vine tree sequence from $T_{1}$ to $T_{7}$ is presented in Fig. F.5.
@@
Returning to model comparison, we consider extremal Hüsler–Reiss graphical models (Engelke et al., 2021; Hentschel et al., 2022). Hentschel et al. (2022) obtained a sparse Hüsler–Reiss graphical model using the EGlearn algorithm (Engelke et al., 2021). The tuning parameter $\rho\geqslant 0$ controls sparsity. The empirical extremal variogram matrix $\widehat{\Gamma}$ is used as edge weight for the minimum spanning tree with smaller elements in $\widehat{\Gamma}$ indicating stronger extremal dependence. Following Engelke et al. (2022), the data set is split in half: a training set for model fitting and a test set for tuning parameter selection. We determine the optimal tuning parameter, $\rho^{*}=0.1$, by evaluating the Hüsler–Reiss log-likelihood across tuning parameter values $\rho$ from the test set. The resulting sparse extremal graph with 148 edges and $\rho^{*}=0.1$ is shown in Fig. LABEL:fig:FlightGraphEGlearn.
@@
We assess the goodness of fit using the entire sub-sample and compare empirical tail dependence coefficients to those obtained from the fitted graphical models as explained in Appendix E.2 in the supplement: the Hüsler–Reiss extremal graphical model for the flight graph in Fig. LABEL:fig:ChiFlightDomain, the truncated X-vine model with $q^{*}=7$ in Fig. LABEL:fig:ChiFlightTruncXvine, and the extremal Hüsler–Reiss graphical model with $\rho^{*}=0.1$ in Fig. LABEL:fig:ChiFlightEGlearn, respectively.
@@
The flexibility of a regular vine structure allows the selection of various (tail) copula families with the lowest AIC. Recall from Section 4.2 that the Hüsler–Reiss model arising from the X-vine specification is constructed using bivariate Hüsler–Reiss tail copula densities in the first tree and bivariate Gaussian copulas in subsequent trees. In the truncated X-vine model, 8 out of 28 edges prefer Hüsler–Reiss models in $T_{1}$, while in trees $T_{j}$ for $j\geqslant 2$, a total of 96 out of 113 edges favour copula families other than the Gaussian one. However, uncertainty in sequential parameter estimation tends to accumulate across tree levels. Consequently, the $\chi$-plot of the truncated X-vine model shows a better fit but more variability. In contrast, the $\chi$-plot of the extremal Hüsler–Reiss graphical model has less variability but seems biased towards higher tail dependence. Fig. F.8 in the supplement presents a model comparison similar to Figs. LABEL:fig:ChiFlightDomainLABEL:fig:ChiFlightEGlearn, but focusing exclusively on the Hüsler–Reiss model and using empirical tail dependence coefficients derived from the empirical extremal variogram matrix.
@@ 10 Discussion
We have opened the door to the construction and the use of extremal dependence models based on regular vine tree sequences. For (ordinary) copulas, the methodology has been extensively developed in the literature since its inception more than two decades ago. Our contribution is to deliver the theoretical and methodological advances required to apply vine machinery in the extreme value context too. The key consists of a version of Sklar’s theorem applied to tail copula densities, together with a telescoping product formula for regular vines.
@@ Acknowledgments
The authors thank Manuel Hentschel for providing the US flight delay data and gratefully acknowledge two reviewers and an associate editor for valuable comments. This work is supported by funds from the Fonds de la Recherche Scientifique – FNRS, Belgium (grant T.0203.21).
Supplementary material for
“X-Vine Models for Multivariate Extremes”
In Section A, we provide detailed illustrations of a number of concepts, results and methods for a particular five-dimensional regular vine sequence. Section B contains the proofs of all results in the paper. For readers more familiar with exponent measure densities than with tail copula densities, a number of theoretical results are reformulated in terms of exponent measure densities in Section C. Section D provides more details on the estimation of the copula parameters in the trees after the first one. Expressions for tail dependence coefficients $\chi_{J}$ (analytical, empirical, and via Monte Carlo) are provided in Section E. Finally, Section F contains additional numerical results, both for the Monte Carlo simulations in Section 8 and the US flight data case study in Section 9.
@@ Appendix A Example in dimension five
In tree $T_{3}$, consider the edge $e=\{a,b\}=\{\{\{1,2\},\{2,3\}\},\{\{2,3\},\{2,4\}\}\}\in E_{3}$, which is also a node in tree $T_{4}$. The proximity condition is satisfied since $|a\cap b|=|\{\{2,3\}\}|=1$: note that $a\cap b$ is indeed a singleton, the unique element of which is the set $\{2,3\}$. The edge $e$ has complete union $A_{e}=\{1,2,3,4\}$, which is the union of the conditioning sets $A_{a}=\left\{{1,2,3}\right\}$ and $A_{b}=\left\{{2,3,4}\right\}$. Further, $e$ has conditioning set $D_{e}=\{2,3\}=A_{a}\cap A_{b}$, and conditioned set $\mathcal{C}_{e}=A_{a}\operatorname{\bigtriangleup}A_{b}=\{1,4\}$, which is itself the union $\mathcal{C}_{e}=\mathcal{C}_{e,a}\cup\mathcal{C}_{e,b}$ of $\mathcal{C}_{e,a}=A_{a}\setminus A_{b}=\left\{{1}\right\}$ and $\mathcal{C}_{e,b}=A_{b}\setminus A_{a}=\left\{{4}\right\}$. Writing this edge in its conditioned form gives better readability: $e=(\mathcal{C}_{e,a},\mathcal{C}_{e,a};D_{e})=(14;23)$. Fig. A.1 shows the regular vine sequence, where the nodes and edges have been labeled using the conditioned forms.
@@
Figure A.1: A graphical representation of the five-dimensional regular vine sequence $\mathcal{V}$. The sub-vine $\mathcal{V}_{f}$ induced by $\mathcal{V}$ on the node set $A_{f}=\{2,4,5\}$ for edge $f=(25;4)$ in $T_{2}$ is shown in blue.
@@
As an illustration of Remark B.4, consider the edge $f=\{\{2,4\},\{4,5\}\}\in E_{2}$ (or $f=(25;4)$ in its conditioned form). The vine $\mathcal{V}_{f}$ induced by $\mathcal{V}$ on the node set $A_{f}=\{2,4,5\}$ consists of two trees, $\mathcal{V}_{f}=(T_{f,1},T_{f,2})$, with edge sets $E_{f,1}=\{\{2,4\},\{4,5\}\}$ and $E_{f,2}=\{\{\{2,4\},\{4,5\}\}\}$; see Fig. A.1. The product formula (B.1) yields the identity
@@
where $C_{1|3;2}(u_{1}|u_{3})=\int_{0}^{u_{1}}c_{13;2}(v,u_{3})\,\mathrm{d}v$ and $R_{1|2}(x_{1}|x_{2})=\int_{0}^{x_{1}}r_{12}(t,x_{2})\,\mathrm{d}t$ and $R_{3|2}(x_{3}|x_{2})=\int_{0}^{x_{3}}r_{23}(x_{2},t)\,\mathrm{d}t$, i.e., in terms of bivariate (tail) copula densities $r_{12}$, $r_{23}$ and $c_{13;2}$ only.
@@
Proofs involving regular vine sequences often operate via a recursive argument in which a node $a\in[d]$ and the edges that contain it are peeled off, leaving a regular vine sequence on $d-1$ elements; see for instance the proof of Theorem 5.2, which relies on Lemma B.3 below. Let us illustrate the latter for the regular vine sequence $\mathcal{V}$ in Figure A.1. The conditioned set of the deepest edge $e=(15;234)$ is $\mathcal{C}_{e}=\left\{{1,5}\right\}$. For each of the latter two nodes and for each level $j\in\left\{{1,\ldots,d-1}\right\}$, there is a unique edge $e_{j}\in E_{j}$ that contains this node in its complete union. For instance, for node $1$ these are the edges $(12)$, $(13;2)$, $(14;23)$ and $(15;234)$; moreover, node $1$ belongs to the conditioned set of each of these edges. Removing node $1$ and these four edges leaves a regular vine sequence on the four remaining elements $\left\{{2,\ldots,5}\right\}$, which is the sub-vine induced by $\mathcal{V}$ on the node set $\left\{{2,\ldots,5}\right\}$.
@@
Steps (2) and (4.a) leave freedom for choice, and this is why there may be multiple permutations $\sigma_{j}$ that satisfy the requirements of Lemma 6.2. For the regular vine tree $\mathcal{V}$ in Fig. A.1, five possible permutations $\sigma_{j}$ for $j\in\left\{{1,\ldots,5}\right\}$ and the corresponding edge sequences $e_{j,1},\ldots,e_{j,4}$ are listed in Table 1. Another choice for $\sigma_{4}$ would have been $(4,2,5,3,1)$, for instance.
@@
Table 1: Five possible permutations $\sigma_{j}$ and the edge sequences $e_{j,1},\ldots,e_{j,4}$ for $j\in\left\{{1,\ldots,5}\right\}$ of Lemma 6.2 for the regular vine tree $\mathcal{V}$ in Figure A.1.
@@
For the regular vine sequence in Fig. A.1, some possible structure matrices $M$ are
@@
We illustrate the recursive nature of the sequential estimation procedure in Section 7.2 for the regular vine sequence in Fig. A.1. For the edge $e=(25;4)$ in $E_{2}$, we consider pseudo-observations